OpenStudy (anonymous):

Rewrite the expression as a trigonometric function of a single angle measure.

OpenStudy (anonymous):

\[\frac{ \tan4 \theta + \tan 2 \theta }{ 1 - \tan 4 \theta + \tan 2 \theta }\]

OpenStudy (anonymous):

hi. do you still need help with this one?

OpenStudy (anonymous):

Yes, I'm so lost on how to solve it

OpenStudy (anonymous):

do you know the tangent addition formula?

OpenStudy (anonymous):

tan(a+b)=(tan a +tan b)/(1-tan a tanb)

OpenStudy (anonymous):

have you seen this formula before?

OpenStudy (anonymous):

now think about the angle a=4theta and the angle b=2theta. does this help?

OpenStudy (anonymous):

Ive seen it but never used it,

OpenStudy (anonymous):

ah. well, your teacher is giving you the opportunity here

OpenStudy (anonymous):

I'm confused on how I solve it with the this formula using the 'theta'

OpenStudy (anonymous):

ok. well, it is more recognition than solving. look at the form of the equation i gave you and the relation you wrote above. write this puppy as the tangent of the sum of the two angles you have been given (4theta and 2theta)

OpenStudy (anonymous):

it is good to ask questions if you are in doubt. :)

OpenStudy (anonymous):

Sorry, If it takes to long to respond, I'm still confused a little so I'm trying to write it down so that would hopefully help me get it

OpenStudy (anonymous):

no problem. so the form of the equation is tan (a+b)=(tan a +tan b)/(1-tan a tan b). substitute for a 4theta and for b 2theta

OpenStudy (anonymous):

So, tan 4theta + tan 2theta / 1- tan 4theta * tan 2.?

OpenStudy (anonymous):

we are interested in the left side. the tan (a+b) part. this condenses the mess you were given.

OpenStudy (anonymous):

So 4 theta + 2 theta.?

OpenStudy (anonymous):

which becomes tan ?

OpenStudy (anonymous):

You would add the 4 & 2 together & get 6 & 6 would be the tan wouldn't it.?

OpenStudy (anonymous):

perfect. yes!! so you tan 6theta, which is the trig function (tan) of a single angle 6theta.

OpenStudy (anonymous):

Alright thank you for help.! :-) I understand this now.!